An experienced bricklayer can construct a small wall in 2 hours. An apprentice can complete the job in 6 hours. Find how long it takes if they work together.
Found 2 solutions by nyc_function, nerdybill:
Answer by nyc_function(2741)

1/2 + 1/6 = 1/x

Solve for x to find your answer.

Answer by nerdybill(7384)

An experienced bricklayer can construct a small wall in 2 hours. An apprentice can complete the job in 6 hours. Find how long it takes if they work together.
.
Let x = time (hours) it takes if they work together
then
x(1/2 + 1/6) = 1
.
Solving for x:
x(1/2 + 1/6) = 1
Multiply both sides by 12:
x(6 + 2) = 12
8x = 12
x = 12/8
x = 3/2
x = 1.5 hours